OpenStudy (anonymous):

okay ppl i dont understand i just needed it done Is (-4, 2) a solution of 3y - 2x = 14?

OpenStudy (anonymous):

Given formula 3y - 2x = 14. To know if (-4,2) is a solution of the equation, just substitute in to see if it fits. 3(2) - 2(-4) = 14 6 + 8 = 14 Since 6+8=14 is true , (-4,2) is a solution of the given equation.

OpenStudy (anonymous):

yes, 3(2) - 2(-4) = 14 6 - -8 =14 6+8 =14